OpenStudy (anonymous):

The Breaker's Manufacturing Company makes 80% of a particular sensor, the Cartin Company makes 15% of them, and the Flutes Company makes the other 5%. The sensors made by Breaker's have a 4% defect rate, the Cartin sensors have a 6% defect rate, and the Flutes sensors have a 9% defect rate. If a sensor is randomly selected from the general population of all sensors, what is the probability that it is defective, given that it was made by Breaker’s? 0.04 0.05 0.15 0.80

OpenStudy (anonymous):

@Agl202 please help me

OpenStudy (anonymous):

@Agl202 it was wrong :(

jimthompson5910 (jim_thompson5910):

Let's say you had 100 sensors 80% are made by Breakers, so 0.80*100 = 80 4% of the 80 are defective, 0.04*80 = 3.2 on average

jimthompson5910 (jim_thompson5910):

divide the number defective in Breakers by the total amount Breakers made 3.2/80 = 0.04

OpenStudy (anonymous):

@jim_thompson5910 ahh thanks so much but I already got it wrong thanks so much though for your help if I see the question again on my retake I will let u know if your right ")